OpenStudy (anonymous):

Express 10^12 miners in examiners. Answer in units of Eminer

OpenStudy (anonymous):

Exa means 10\(^{18}\), which means you have to divide your number of miners by 10\(^{18}\)L \[\#Examiners=\frac{10^{12}}{10^{18}}\]When you're dividing powers with the same base, just subtract the powers: \[\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (nilankshi):

hmmm..................